1. An analysis of monthly wages paid to the workers in two company X and Y gives the following results:  

 

 

                                                                      Firm X                Firm Y 

Number of worker:                            45                    50 

Average monthly wages in $:            100                  120  

Variance of distribution of wage:       49                     36             

  1. Which company, X or Y, has a large wage bill? 
  2. In which company, X or Y is there grater variability in wage? 
  3. Calculate combined mean wage of company X and Y
1
Expert's answer
2021-01-11T18:41:42-0500

1. Company Y has a larger wage bill.

2. In the Company X there is a larger variability in wages.

3. Combined mean wage: "\frac{45*100+50*120}{100+120}=\frac{10500}{220}=47.73."